# Spectral Thresholds for Identifiability and Stability: Finite-Sample Phase Transitions in High-Dimensional Learning

### v1

@@ 1 Introduction
In modern high-dimensional learning, models often appear reliable up to a point, only to collapse abruptly once the sample size falls below a critical level. This is evident in double descent in overparameterized neural networks [Belkin et al., 2019] and in high-dimensional regression, where estimation becomes unreliable once $n$ is on the order of $d$. Existing frameworks—such as Fisher consistency, restricted eigenvalue conditions, or information-theoretic bounds—offer only asymptotic guarantees or loose sufficient criteria, leaving these sharp transitions unexplained.
@@
Main result. Our central theorem establishes a Fisher spectral threshold: identifiability requires the minimal eigenvalue of the empirical Fisher information to exceed an explicit $O(\sqrt{d/n})$ bound. Unlike asymptotic Fisher consistency [Le Cam, 1970], this threshold is both necessary and finite-sample: above it, parameters concentrate reliably; below it, estimation fails due to Fisher spectrum degeneracy, as weak eigendirections become indistinguishable under finite-sample noise. This refines prior phase-transition analyses in spiked models [Baik et al., 2005] and double descent [Belkin et al., 2019], yielding a sharp non-asymptotic boundary for when learning remains possible. Detailed proofs, including the PL inequality, are in Section 4, building on assumptions in Section 3.
@@ 2 Related Work
Classical asymptotic statistics links identifiability to Fisher information, via local asymptotic normality and the Cramér–Rao inequality [Le Cam, 1970, van der Vaart, 1998]. In high-dimensional settings, conditions such as restricted eigenvalue and restricted strong convexity [Bickel et al., 2009, Negahban et al., 2012] yield sufficient guarantees, while modern analyses reveal sharp feasibility boundaries for specific MLEs [Sur and Candès, 2019]. However, these results are either asymptotic, provide only loose sufficient criteria, or remain tied to particular model classes, leaving open whether there exists a verifiable necessary law that dictates when identifiability must collapse. Our contribution addresses this gap by providing a general necessary finite-sample spectral law, reframing Fisher information as a concrete non-asymptotic criterion for identifiability.
@@
Phase-transition phenomena are central in high-dimensional inference: the BBP transition in spiked models [Baik et al., 2005], sparse PCA [Lesieur et al., 2015], and multi-index models [Defilippis et al., 2025]. In machine learning, related instabilities appear through the “double descent” phenomenon [Belkin et al., 2019] and analyses of Fisher information spectra in deep networks [Pennington and Worah, 2018, Karakida et al., 2019]. These works demonstrate that spectral degeneracies often coincide with instability, but their conclusions remain either tied to specific models or descriptive in nature, and therefore stop short of providing general, verifiable necessary thresholds. Our results sharpen these insights by establishing an explicit spectral boundary that marks the onset of stability failure, connecting Fisher spectrum degeneracy directly to finite-sample identifiability as a necessary law.
@@
Within learning theory, stability and generalization have been studied through uniform stability [Bousquet and Elisseeff, 2002, Hardt et al., 2016], PAC-Bayesian analysis [McAllester, 1999, Dziugaite and Roy, 2017], and information-theoretic approaches [Xu and Raginsky, 2017]. These frameworks largely provide sufficient guarantees, ensuring generalization when stability holds, but do not characterize when stability must necessarily fail. Lower bounds, such as those of Feldman and Vondrak [2018], highlight the inherent limits of uniform stability, but remain tied to specific algorithmic assumptions. Our contribution complements this line by establishing a spectral lower bound on stability: an algorithm-independent criterion that becomes binding once Fisher curvature falls below sample-level fluctuations, thereby marking a fundamental and verifiable impossibility frontier for learning algorithms.
@@ 3 Preliminaries
    Spectral perturbation (Weyl). [Stewart and Sun, 1990] $\lambda_{\min}(\widehat{\Gamma})\geq\lambda_{\min}(\Gamma)-\Lambda^{\ast}$, where $\Lambda^{\ast}=C_{0}\sigma_{\mathrm{eff}}\sqrt{\tfrac{d+\log(1/\delta)}{n}}$.
@@ 4.1 Fisher Spectral Threshold (Theorem 1)
yielding linear convergence of gradient descent [Karimi et al., 2016, Polyak, 1963]. Conversely, if $\lambda_{\min}(\Gamma)\leq\tfrac{1}{2}\Lambda^{\ast}$, then indistinguishable local alternatives exist (via Le Cam) [Le Cam and Yang, 2000], so identifiability collapses and no uniform PL inequality can hold.
@@ 4.4 Preconditioning Robustness (Proposition 5)
Under whitening with $(1-\alpha)\Sigma\preceq\widehat{\Sigma}\preceq(1+\alpha)\Sigma$ [Bishop, 2006, Ba et al., 2016], the minimal eigenvalue is perturbed only by $(1\pm\alpha)^{-1}$ constants in the $\Sigma$-geometry.
@@ 5 Experiments: Spectral Phase Transitions in Practice
Finally, Corollary 4.4 states that stability can be certified using only a finite set of directions, with error controlled by the angle penalty. The tracked Rayleigh minimum $\phi_{K}$ consistently upper-bounds the true $\lambda_{\min}$ and converges to it over time, while the residual gap decays exactly as $\Delta_{B}\sin^{2}\vartheta$ [Davis and Kahan, 1970]. This establishes finite-direction monitoring as a practical tool: stability can be certified online without full spectral computation, with discrepancy precisely governed by the angle term.
@@ 6 Discussion
The results are local around $\theta^{\ast}$; extending to global nonconvex landscapes remains open. They rely on robust Fisher concentration, which heavy-tailed or adversarial settings may violate. Regularization by Fisher floors introduces computational overhead, though randomized sketching (e.g. Hutch++) provides scalable approximations. [Meyer et al., 2021] These caveats also highlight future research: global analysis, robust concentration, and efficient spectral monitoring.

### v2

@@ 1 Introduction
In modern high-dimensional learning, models often appear reliable up to a point, only to collapse abruptly once the sample size falls below a critical level. This is evident in double descent in overparameterized neural networks [belkin2019reconciling] and in high-dimensional regression, where estimation becomes unreliable once $n$ is on the order of $d$. Existing frameworks—such as Fisher consistency, restricted eigenvalue conditions, or information-theoretic bounds—offer only asymptotic guarantees or loose sufficient criteria, leaving these sharp transitions unexplained.
@@
Main result. Our central theorem establishes a Fisher spectral threshold: identifiability requires the minimal eigenvalue of the empirical Fisher information to exceed an explicit $O(\sqrt{d/n})$ bound. Unlike asymptotic Fisher consistency [lecam1970asymptotic], this threshold is both necessary and finite-sample: above it, parameters concentrate reliably; below it, estimation fails due to Fisher spectrum degeneracy, as weak eigendirections become indistinguishable under finite-sample noise. This refines prior phase-transition analyses in spiked models [baik2005phase] and double descent [belkin2019reconciling], yielding a sharp non-asymptotic boundary for when learning remains possible. Detailed proofs, including the PL inequality, are in Section 4, building on assumptions in Section 3.
@@ 2 Related Work
Classical asymptotic statistics links identifiability to Fisher information, via local asymptotic normality and the Cramér–Rao inequality [lecam1970asymptotic, vandervaart1998asymptotic]. In high-dimensional settings, conditions such as restricted eigenvalue and restricted strong convexity [bickel2009simultaneous, negahban2012unified] yield sufficient guarantees, while modern analyses reveal sharp feasibility boundaries for specific MLEs [sur2019modern]. However, these results are either asymptotic, provide only loose sufficient criteria, or remain tied to particular model classes, leaving open whether there exists a verifiable necessary law that dictates when identifiability must collapse. Our contribution addresses this gap by providing a general necessary finite-sample spectral law, reframing Fisher information as a concrete non-asymptotic criterion for identifiability.
@@
Phase-transition phenomena are central in high-dimensional inference: the BBP transition in spiked models [baik2005phase], sparse PCA [lesieur2015phase], and multi-index models [defilippis2025optimal]. In machine learning, related instabilities appear through the “double descent” phenomenon [belkin2019reconciling] and analyses of Fisher information spectra in deep networks [pennington2018spectrum, karakida2019universal]. These works demonstrate that spectral degeneracies often coincide with instability, but their conclusions remain either tied to specific models or descriptive in nature, and therefore stop short of providing general, verifiable necessary thresholds. Our results sharpen these insights by establishing an explicit spectral boundary that marks the onset of stability failure, connecting Fisher spectrum degeneracy directly to finite-sample identifiability as a necessary law.
@@
Within learning theory, stability and generalization have been studied through uniform stability [bousquet2002stability, hardt2016train], PAC-Bayesian analysis [mcallester1999some, dziugaite2017computing], and information-theoretic approaches [xu2017information]. These frameworks largely provide sufficient guarantees, ensuring generalization when stability holds, but do not characterize when stability must necessarily fail. Lower bounds, such as those of feldman2018generalization, highlight the inherent limits of uniform stability, but remain tied to specific algorithmic assumptions. Our contribution complements this line by establishing a spectral lower bound on stability: an algorithm-independent criterion that becomes binding once Fisher curvature falls below sample-level fluctuations, thereby marking a fundamental and verifiable impossibility frontier for learning algorithms.
@@ 3 Preliminaries
    Spectral perturbation (Weyl). [stewart1990matrix] $\lambda_{\min}(\widehat{\Gamma})\geq\lambda_{\min}(\Gamma)-\Lambda^{\ast}$, where $\Lambda^{\ast}=C_{0}\sigma_{\mathrm{eff}}\sqrt{\tfrac{d+\log(1/\delta)}{n}}$.
@@ 4.1 Fisher Spectral Threshold (Theorem 1)
yielding linear convergence of gradient descent [karimi2016linear, polyak1963gradient]. Conversely, if $\lambda_{\min}(\Gamma)\leq\tfrac{1}{2}\Lambda^{\ast}$, then indistinguishable local alternatives exist (via Le Cam) [lecam2000asymptotics], so identifiability collapses and no uniform PL inequality can hold.
@@ 4.4 Preconditioning Robustness (Proposition 5)
Under whitening with $(1-\alpha)\Sigma\preceq\widehat{\Sigma}\preceq(1+\alpha)\Sigma$ [bishop2006pattern, ba2016layernorm], the minimal eigenvalue is perturbed only by $(1\pm\alpha)^{-1}$ constants in the $\Sigma$-geometry.
@@ 5 Experiments: Spectral Phase Transitions in Practice
Finally, Corollary 4.4 states that stability can be certified using only a finite set of directions, with error controlled by the angle penalty. The tracked Rayleigh minimum $\phi_{K}$ consistently upper-bounds the true $\lambda_{\min}$ and converges to it over time, while the residual gap decays exactly as $\Delta_{B}\sin^{2}\vartheta$ [daviskahan1970rotation]. This establishes finite-direction monitoring as a practical tool: stability can be certified online without full spectral computation, with discrepancy precisely governed by the angle term.
@@ 6 Discussion
The results are local around $\theta^{\ast}$; extending to global nonconvex landscapes remains open. They rely on robust Fisher concentration, which heavy-tailed or adversarial settings may violate. Regularization by Fisher floors introduces computational overhead, though randomized sketching (e.g. Hutch++) provides scalable approximations. [hutchpp2021] These caveats also highlight future research: global analysis, robust concentration, and efficient spectral monitoring.